OpenStudy (anonymous):

what is equivalent to the expression below. select all that apply. 3^4*3^x A. 3^(4-x) B. 81*(3^x) C. (3*x)^4 D. 3^(4x) E. 3^(4+x) F. 9^(4x)

OpenStudy (blurbendy):

actually

OpenStudy (blurbendy):

since the base of the exponents are the same, do you remember what you can do with that fact?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

so B is an answer

OpenStudy (anonymous):

then.... what else

OpenStudy (blurbendy):

if you have two bases that are the same: a^2 + a^2, then you can add the exponents to get: a^(2+2)

OpenStudy (blurbendy):

so B is an answer and so is E

OpenStudy (anonymous):

oops I forgot-sorry it was on the very bottom of my notes

OpenStudy (anonymous):

so yes E is also an answer

OpenStudy (anonymous):

is that it though?

OpenStudy (blurbendy):

yes, none of the other ones work

OpenStudy (anonymous):

omg! did really bad on a few of thse questions so can u help me w/ like a few more questions pls??

OpenStudy (blurbendy):

ok

OpenStudy (anonymous):

ok i'll post it now

OpenStudy (anonymous):

which are equivalent below,. select all that apply. 9^x A. 9*(9^(x-1)) B. (36/4)^x C. 9*(9^(x+1)) D. (36^x)/4 E. 36^x/4^x F. x^5

OpenStudy (anonymous):

ok so I know that B & E will work

OpenStudy (anonymous):

but im not sure abt the others

OpenStudy (blurbendy):

A works as well

OpenStudy (anonymous):

ok so A, B & E ?

OpenStudy (anonymous):

but how is A though?

OpenStudy (blurbendy):

9 * 9^(x - 1) = 9 * 1/9 * 9^x

OpenStudy (blurbendy):

9 and 1/9 cancel out to equal 1, and 1 * 9^x = 9^x

OpenStudy (blurbendy):

9^(-1) = 1/9

OpenStudy (anonymous):

ok the next one is the same type of problem w/ same intructions

OpenStudy (blurbendy):

ok

OpenStudy (anonymous):

question: 2^5*2^x A. 2^(5x) B. 4^(5x) C. 2^(5-x) D. (2*x)^5 E. 2^(5+x) F. 32*(2^x)

OpenStudy (blurbendy):

2^5 is 32 so F will work. 2^(5+x) will also work (same bases mead we can add the exponents. none of the others work

OpenStudy (anonymous):

yeah I had F

OpenStudy (anonymous):

wait E also works too

OpenStudy (anonymous):

so E & F

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

next question: 16^x A. 4*(4^x) B. 4*(4^2x) C. 4^2*4^x D. 4^(2x) E. 4^x*4^x F. (4*4)^x

OpenStudy (anonymous):

I know E & F works

OpenStudy (anonymous):

is that it? that's all I could find

OpenStudy (blurbendy):

d works

OpenStudy (anonymous):

4^(2x)?? how?

OpenStudy (blurbendy):

4^(2x) = 4^2*4^x which is the same as E

OpenStudy (blurbendy):

wait no it's not

OpenStudy (anonymous):

yea.... cuz then they'd have to add together not multiply

OpenStudy (blurbendy):

wait

OpenStudy (anonymous):

ok

OpenStudy (blurbendy):

it might be a good idea to get a second opinion because im not sure actually

OpenStudy (anonymous):

ok..

OpenStudy (blurbendy):

just post the question

OpenStudy (anonymous):

i'm gonna close this question and put the one we were discussing

OpenStudy (blurbendy):

kk